Scientific Notation in problem solving.
Named in tribute to Anders Angstrom, a Swedish physicist who measured light waves, 1 A (read "one Angstrom") equals 10^-10 meter. One parsec is about 3.26 light years, and one light year equals 9.46 x 10^15 meters.
How many Angstroms are in one parsec?
My answer was 3.08 x 10^6. The books answer was 3.08 x 10^26. If I'm right, let me know. If I'm wrong, show me how it's done.

Answer by rothauserc(4718)

(3 * (9.46 * 10^15)) / 10^(-10) = 2.8 x 10^26 approx 3 x 10^26 Angstroms
note that 1 / 10^(-10) = 10^10